A mixture of mthane and ethane in the mole ratio x:y has a mean molecular mass equal to 20 . What would be the mean molecular mass of a mixture of tha same gases present in the ratio y:x?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Shivi
 

Molar ratio of methane : ethene = a:b
Average molecular mass of the mixture = 16a + 28b / a+b

Now,

16a + 28b / a+b  = 20

On solving,

a/b = 2/1

On reversing b/a = 1/2

so, 1:2 = (16 x1 + 28x 2)/3 = 24

 

Regards

Arun (askIITians forum expert)